# ManiFeel: Benchmarking and Understanding Visuotactile Manipulation Policy Learning

### v1

@@ A.1 Imitation Learning
A.2 Visuotactile Manipulation
Tactile feedback has been widely explored for manipulation tasks involving rigid [34, 35, 36, 37], deformable [38, 39, 40], and articulated objects [41]. A common strategy is to extract low-level state estimations, such as contact areas or object pose, from tactile signals to guide model-based controllers. Moreover, some approaches integrate tactile signals into model predictive control frameworks [42], or into task-specific solutions like page flipping [43], grasping [44], robotic pushing [45]. Finally, learning-based tactile control frameworks have emerged to improve adaptability. For instance, neural networks have been used to predict grasp stability or success [46, 47, 48], and to evaluate candidate actions by sampling-based method [49, 50, 51, 52].
@@ A.2 Visuotactile Manipulation
A.3 Tactile Representation Learning

### v2

@@ II-A Imitation Learning
II-B Visuotactile Manipulation
Tactile feedback has been widely explored for manipulation tasks involving rigid [22, 28, 42, 12], deformable [41, 44, 2, 68], and articulated objects [66]. A common strategy is to extract low-level state estimations, such as contact areas or object pose, from tactile signals to guide model-based controllers. Moreover, some approaches integrate tactile signals into model predictive control frameworks [47, 33], or into task-specific solutions like page flipping [67], grasping [52], robotic pushing [32, 34]. Finally, learning-based tactile control frameworks have emerged to improve adaptability. For instance, neural networks have been used to predict grasp stability or success [8, 43, 27], and to evaluate candidate actions by sampling-based method [7, 23, 13, 18].
@@ II-B Visuotactile Manipulation
II-C Tactile Representation Learning